# Generic groups and the weak amalgamation property

### v2

@@ 3.2 Generics over rationals
We now give an application of Theorem 3.2 for actions on $\mathbb{Q}$ (and thus on the real line $\mathbb{R}$). Consider Scenario II of Section 1.2, where $M_{0}=(\mathbb{Q},<)$. We remind the reader that this structure is $\omega$-categorical, admits elimination of quantifiers and its theory is decidable. Furthermore, these properties still hold for any expansion $(\mathbb{Q},<,\bar{a})$ by a tuple $\bar{a}\in\mathbb{Q}$. Note that the theory of this expansion is determined by the size of $\bar{a}$.
Now define $M$ to be $M_{0}$ with the additional sort $\mathsf{Gp}$ and the constant symbol $1$ interpreted by $1\in\omega$. The symbols $\bar{r}$ are $\cdot$, -1 and $\mathsf{ac}$ for an action $M_{0}\times\mathsf{Gp}\rightarrow M_{0}$ by automorphism.
The theory $T$ contains the universal axioms of groups on the sort $\mathsf{Gp}$ with the unit $1$ and the axioms for an action by automorphisms. The space $\mathcal{X}_{M,T}$ is the logic space of all countable expansions of $M$ where the group structure is defined on $\mathsf{Gp}$, with a fixed unit $1$ and an action $\mathsf{ac}$ We denote it by $\mathcal{X}_{(\mathbb{Q},<)}$.
@@
As in Corollary 3.3 we use the construction of A. Darbinyan of a finitely presented left orderable group $G=\langle\bar{c}|\mathcal{R}(\bar{c})\rangle$ with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that any homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$ has undecidable word problem (see Proposition 5.2.13 of [6]).
The group $G$ acts on $(\mathbb{Q},<)$ by automorphisms. Indeed, take the ordering $<$ on $G$ so that $G$ acts on it from the left. If $<$ is dense we identify it with $\mathbb{Q}$. If $<$ is discrete then embed it into $(\mathbb{Q},<)$. Its convex hull in $\mathbb{Q}$ can be identified with $(\mathbb{Q},<)$ again (it is easy to see that there is no maximal element in $(G,<)$). As a result we decompose $(\mathbb{Q},<)$ into a family of intervals with endpoints in $G$. All of them can be identified with the interval $(1,g_{0})$, where $g_{0}$ is the least positive element of $G$. Now extend the action of $G$ on $G$ (by left multiplication) to the rest of $\mathbb{Q}$ using this identification.
Now by Theorem 3.2 we see that $\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c})\not=w_{2}(\bar{c})\}$ defines a diagram without extensions witnessing WAP. $\Box$
@@ 4 Semigroups and rings
The approach of Section 3 can be easily generalized to other structures studied in algebra. In this section we consider how it works in the cases of semigroups and rings.
In the cases of semigroups we modify Scenario I of Section 1.2 so that the structure $M$ is just $\omega$ and $L$ is empty. Let $\bar{r}$ consist of the binary function of multiplication   $\cdot$ . Let $T$ be the universal theory of semigroups. Under this convention each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ and a description of a partial function for $\cdot$ on $\bar{c}$ (the associativity axiom for $\cdot$ is also added), i.e. a partial semigroup. The space $\mathcal{X}_{M,T}$ is the logic space of all countable semigroups. We denote it by $\mathcal{S}$.
The case of rings is similar. We now assume that $\bar{r}$ consist of three binary functions of addition   $+$   , subtraction   --   and multiplication   $\cdot$ . We also include the constant symbol 00 into $\bar{r}$. Let $T$ be the universal theory of associative rings (without unit). Thus each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ and a description of three partial functions for $+$, -- and $\cdot$ on $\bar{c}$ (with 00 and axioms of a ring). The space $\mathcal{X}_{M,T}$ corresponding to $T$ will be denoted by $\mathsf{Ring}$.
Completeness and Kuznetsov’s theorem for semigroups and rings work as follows. Firstly define t-isolation.
@@
Let $\bar{c}$ be a subtuple of $\bar{c}^{\prime}$ and $D(\bar{c}^{\prime})\in\mathcal{B}_{T}$. We say that $D(\bar{c}^{\prime})$ is t-isolating for $\bar{c}$ if in every member of $[D(\bar{c}_{1})]$ the semigroup (resp. ring) $\langle\bar{c}\rangle$ satisfies the same semigroup (resp. ring) equalities of the form $w_{1}(\bar{c}^{\prime})=w_{2}(\bar{c}^{\prime\prime})$ with $\bar{c}^{\prime},\bar{c}^{\prime\prime}\subseteq\bar{c}$.
Proposition 3.1 can be adapted as follows.
Proposition 4.2
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ and an extension $D^{\prime}(\bar{c}^{\prime})\in\mathcal{B}_{T}$ is t-isolating for $\bar{c}$. Let $(M,{\bf\cdot})$ (resp. $(M,+,0,-,\cdot)$) $\in[D^{\prime}(\bar{c}^{\prime})]$. Assume that $D^{\prime}(\bar{c}^{\prime})$ is a computably enumerable set. Then in $(M,{\bf\cdot})$ (resp. in $(M,+,-,0,\cdot)$), the elements of $\bar{c}$ generate a recursively presented semigroup (resp. ring) with decidable word problem.
The proof of this proposition coincides with the proof of Proposition 3.1. The only changes concern the form of equations and inequations appearing there. In the semigroup case we should replace expressions $w=(\not=)1$ by $w_{1}=(\not=)w_{2}$ where $w_{1}$ and $w_{2}$ are words over $\bar{c}$. In the ring case we consider these expressions again, but now $w_{1}$ and $w_{2}$ will be polynomials on $\bar{c}$.
Theorem 3.2 also holds in the cases of semigroups and rings. The proof is the same (with changes as above).
Theorem 4.3
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ satisfies the property that every extension $D^{\prime}(\bar{c}^{\prime})$ is computably enumerable and for every $(M,{\bf\cdot})$ (resp. $(M,+,0,-,\cdot)$) $\in[D(\bar{c})]$ the semigroup (resp. ring) $\langle\bar{c}\rangle$ (defined in $(M,{\bf\cdot})$ or $(M,+,0,-,\cdot)$) has undecidable word problem.
@@
Theorem 4.3 is applied to enumerated semigroups as follows.
@@
Proof. It is obvious that all members of $\mathcal{B}_{T}$ are computably enumerable. We now repeat the argument of Corollary 3.3. Let $G=\langle\bar{c}\,|\,\mathcal{R}(\bar{c})\rangle$ be a finitely presented left orderable group with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that any homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$, has undecidable word problem (see Proposition 5.2.13 of [6]). This group can be converted into a finitely presented semigroup as follows. Let $n=|\bar{c}|$. For every $c_{i}\in\bar{c}$ we introduce the letter $c_{n+i}$. The group $G$ has a semigroup presentation for generators $\bar{c},1,c_{n+1},\ldots,c_{2n}$ as follows. We start with relations saying that $1$ is the unit: $c_{i}\cdot 1=1\cdot c_{i}=c_{i}$, $1\leq i\leq 2n$ and $1\cdot 1=1$. Now add all relations of the form $c_{i}\cdot c_{n+i}=c_{n+i}\cdot c_{i}=1$, $1\leq i\leq n$. Every word from $\mathcal{R}(\bar{c})$ can be rewritten as a semigroup word by replacing each $c^{-1}_{i}$ by $c_{n+i}$. We extend the family of relations by all equalities $w(\ldots,c_{i},\ldots c_{n+i},\ldots)=1$ for all words rewriting words from $\mathcal{R}(\bar{c})$. It is easy to see that the semigroup with this presentation is isomorphic to $G$. Let $w^{\prime}_{1}$ and $w^{\prime}_{2}$ be obtained from $w_{1}(\bar{c})$ and $w_{2}(\bar{c})$ using the rewriting procedure described above for $\mathcal{R}(\bar{c})$. Applying Theorem 4.3 we see that the semigroup presentation of $G$ together with $w^{\prime}_{1}\not=w^{\prime}_{2}$ is a diagram without extensions witnessing WAP. This follows from the argument of Corollary 3.3 together with the observation that every semigroup homomorphism from $G$ is a group homomorphism. $\Box$
@@
Let $\pi_{0}$ be the quotient map from $A_{0}$ to $A$. Applying the construction of Belyaev from [3] (described in the beginning of our proof) we computably embed $A$ into a finitely generated ring $R$, which is finitely presented in $\mathsf{Ring}$. Denote this embedding by $\phi$. Set $F=\{\phi(a_{1}),\phi(a_{2})\}\subset R$. Repeating the argument of Theorem 3.4 it is easy to verify that, for any surjective homomorphism, say $\pi$ from $R$ to a ring $R^{\prime}$ that is injective when restricted to $F$, we have that $R^{\prime}$ has undecidable problem of equality of terms.

### v3

@@ 3.2 Generics over rationals
We now give an application of Theorem 3.2 to actions on $\mathbb{Q}$ (and thus on the real line $\mathbb{R}$). Consider Scenario II of Section 1.2, where $M_{0}=(\mathbb{Q},<)$. We remind the reader that this structure is $\omega$-categorical, admits elimination of quantifiers and its theory is decidable. Furthermore, these properties still hold for any expansion $(\mathbb{Q},<,\bar{a})$ by a tuple $\bar{a}\in\mathbb{Q}$. Note that the theory of this expansion is determined by the size of $\bar{a}$.
Now define $M$ to be $M_{0}$ with the additional sort $\mathsf{Gp}$ and the constant symbol $1$ interpreted by $1\in\omega$. The symbols $\bar{r}$ are $\cdot$, -1 and $\mathsf{ac}$, for an action $M_{0}\times\mathsf{Gp}\rightarrow M_{0}$ by automorphisms.
The theory $T$ contains the universal axioms of groups on the sort $\mathsf{Gp}$ with the unit $1$ and the axioms for an action by automorphisms (i.e. << is preserved). The space $\mathcal{X}_{M,T}$ is the logic space of all countable expansions of $M$ where the group structure is defined on $\mathsf{Gp}$, with a fixed unit $1$ and an action $\mathsf{ac}$ . We denote it by $\mathcal{X}_{(\mathbb{Q},<)}$.
@@
As in Corollary 3.5 we use the construction of A. Darbinyan of a finitely presented left orderable group $G=\langle\bar{c}|\mathcal{R}(\bar{c})\rangle$ with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that any homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$, has undecidable word problem (see Proposition 5.2.13 of [8]).
The group $G$ acts on $(\mathbb{Q},<)$ by automorphisms. Indeed, take the ordering << on $G$ so that $G$ acts on it from the left. If << is dense we identify it with $\mathbb{Q}$. If << is discrete then embed it into $(\mathbb{Q},<)$. Its convex hull in $\mathbb{Q}$ can be identified with $(\mathbb{Q},<)$ again (it is easy to see that there is no maximal element in $(G,<)$). As a result we decompose $(\mathbb{Q},<)$ into a family of intervals with endpoints in $G$. All of them can be identified with the interval $(1,g_{0})$, where $g_{0}$ is the least element of $G$ greater than $1$. Extend the action of $G$ on $G$ (by left multiplication) to the rest of $\mathbb{Q}$ using this identification.
Now by Corollary 3.3 we see that $\mathcal{R}(\bar{c})\cup\{w_{1}(\bar{c})\not=w_{2}(\bar{c})\}$ defines a diagram without extensions witnessing WAP. Applying Theorem 1.2 from [9] we get the result. $\Box$
This result fits to the topic of generic representations of countable groups in Polish groups, for example see [5]. A typical situation studied in these investigations is as follows. Take a countable group $\Gamma$ and a countable first order structure $M$. Now let $\mathsf{Rep}(\Gamma,\mathsf{Aut}(M))$ be the space of all representations of $\Gamma$ in $\mathsf{Aut}(M)$. It is considered under the natural Polish topology (from $\mathsf{Aut}(M)^{\Gamma}$) and the conjugacy action of $\mathsf{Aut}(M)$. The notion of a generic representation in $\mathsf{Rep}(\Gamma,\mathsf{Aut}(M))$ is defined in the standard way, see Introduction in [5]. Note that every representation of $\Gamma$ can be viewed as an expansion $(M,g_{\gamma})_{\gamma\in\Gamma}$ where each $g_{\gamma}$ is the automorphism corresponding to $\gamma$. Thus, under Scenario II of Section 1.2, $\mathsf{Rep}(\Gamma,\mathsf{Aut}(M))$ forms the subset of $\mathcal{X}_{M,T}$ where the group structure of $\mathsf{Gp}$ is fixed (and coincides with $\Gamma$). It is well known that $\mathsf{Rep}(\mathbb{Z},\mathsf{Aut}(\mathbb{Q},<))$ has a generic representation (see [16]), but $\mathsf{Rep}(\mathbb{F}_{2},\mathsf{Aut}(\mathbb{Q},<))$ does not have it (an unpublished theorem of I.Hodkinson). Corollary 3.7 naturally complements these facts. Furthermore, it seems plausible that the fact concerning $\mathbb{F}_{2}$ can be useful for an alternative proof of it.
@@ 4 Semigroups and rings
The approach of Section 3 can be applied to other structures studied in algebra. In this section we consider how it works in the cases of semigroups and rings.
In the cases of semigroups we modify Scenario I of Section 1.2 so that the structure $M$ is just $\omega$ and $L$ is empty (i.e. in terms of Section 1.2 the sort $\mathsf{Gp}$ coincides with $M$). Let $\bar{r}$ consist of the binary function of multiplication   $\cdot$. Let $T$ be the universal theory of semigroups. Under this convention each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ and a description of a partial function for $\cdot$ on $\bar{c}$ (the associativity axiom for $\cdot$ is also added), i.e. a partial semigroup. The space $\mathcal{X}_{M,T}$ is the logic space of all countable semigroups. We denote it by $\mathcal{S}$.
The case of rings is similar. We now assume that $\bar{r}$ consist of three binary functions of addition   ++   , subtraction   -   and multiplication   $\cdot$ . We also include the constant symbol 0 into $\bar{r}$. Let $T$ be the universal theory of associative rings (without unit). Thus, each element of $\mathcal{B}_{T}$ consists of a tuple $\bar{c}\subset\omega$ and a description of three partial functions for ++, - and $\cdot$ on $\bar{c}$ (with 0 and axioms of a ring). The space $\mathcal{X}_{M,T}$ corresponding to $T$ will be denoted by $\mathsf{Ring}$.
Completeness and Kuznetsov’s theorem for semigroups and rings work as follows. It is obvious that in both cases Definition 2.1 is equivalent to the following one.
@@
Let $\bar{c}$ be a subtuple of $\bar{c}_{1}$ and $D(\bar{c}_{1})\in\mathcal{B}_{T}$. We say that $D(\bar{c}_{1})$ is t-isolating for $\bar{c}$ if in every member of $D(\bar{c}_{1})]$ the subsemigroup (resp. subring) generated by $\bar{c}$ in $M$ satisfies the same semigroup (resp. ring) equalities of the form $w_{1}(\bar{c}^{\prime})=w_{2}(\bar{c}^{\prime\prime})$ with $\bar{c}^{\prime},\bar{c}^{\prime\prime}\subseteq\bar{c}$.
The results of Section 3 have natural counterparts for semigroups and rings. For example, the following statement is a counterpart of Corollary 3.3.
Corollary 4.2
Assume that $D(\bar{c})\in\mathcal{B}_{T}$ satisfies the property that for every $(M,{\bf\cdot})$ (resp. $(M,+,0,-,\cdot)$) $\in[D(\bar{c})]$ the subsemigroup (resp. subring) generated by $\bar{c}$ in $(M,{\bf\cdot})$ (resp. $(M,+,0,-,\cdot)$) has undecidable word problem.
@@
The easy form of this corollary follows from the fact that $T$ is computably enumerable in it, and the language of $M$ is empty. Corollary 4.2 is applied to enumerated semigroups as follows.
@@
Proof. We repeat the argument of Corollary 3.5. Let $G=\langle\bar{c}\,|\,\mathcal{R}(\bar{c})\rangle$ be a finitely presented left orderable group with two elements $w_{1}(\bar{c}),w_{2}(\bar{c})\in G$ such that any homomorphic image of $G$ where $w_{1}(\bar{c})\not=w_{2}(\bar{c})$, has undecidable word problem (see Proposition 5.2.13 of [8]). This group can be converted into a finitely presented semigroup as follows. Let $n=|\bar{c}|$. For every $c_{i}\in\bar{c}$ we introduce the letter $c_{n+i}$. The group $G$ has a semigroup presentation for generators $\bar{c},1,c_{n+1},\ldots,c_{2n}$ as follows. We start with relations saying that $1$ is the unit: $c_{i}\cdot 1=1\cdot c_{i}=c_{i}$, $1\leq i\leq 2n$ and $1\cdot 1=1$. Now, add all relations of the form $c_{i}\cdot c_{n+i}=c_{n+i}\cdot c_{i}=1$, $1\leq i\leq n$. Every word from $\mathcal{R}(\bar{c})$ can be rewritten as a semigroup word by replacing each $c^{-1}_{i}$ by $c_{n+i}$. We extend the family of relations by all equalities $w(\ldots,c_{i},\ldots c_{n+i},\ldots)=1$ for all words rewriting ones from $\mathcal{R}(\bar{c})$. It is easy to see that the semigroup with this presentation is isomorphic to $G$. Let $w^{\prime}_{1}$ and $w^{\prime}_{2}$ be obtained from $w_{1}(\bar{c})$ and $w_{2}(\bar{c})$ using the rewriting procedure described above for $\mathcal{R}(\bar{c})$. Applying Corollary 4.2 we see that the semigroup presentation of $G$ together with $w^{\prime}_{1}\not=w^{\prime}_{2}$ is a diagram without extensions witnessing WAP. This follows from the argument of Corollary 3.5 together with the observation that every semigroup homomorphism from $G$ is a group homomorphism. The rest is clear. $\Box$
@@
Let $\pi_{0}$ be the quotient map from $A_{0}$ to $A$. Applying the construction of Belyaev from [3] (described in the beginning of our proof) we computably embed $A$ into a finitely generated ring $R$, which is finitely presented in $\mathsf{Ring}$. Denote this embedding by $\phi$. Set $F=\{\phi(a_{1}),\phi(a_{2})\}\subset R$. Repeating the argument of Corollary 3.6 it is easy to verify that, for any surjective homomorphism, say $\pi$ from $R$ to a ring $R^{\prime}$ that is injective when restricted to $F$, we have that $R^{\prime}$ has undecidable problem of equality of terms.